# Spectral Convolutional Conditional Neural Processes

### v4

@@ B.3.1 Model Architectures
Consistent with the Predator–Prey experiments described in Section B.2.1, the model architectures follow those introduced in Section B.1.1. For the discretization of the functional embedding in both ConvCNP and SConvCNP, we use a margin of 0.5 and a resolution of 48 grid points per unit length.
B.3.2 Training and Data Description
We use the California traffic-flow dataset from LargeST [liu2023largest], a large-scale benchmark for traffic forecasting available at https://www.kaggle.com/datasets/liuxu77/largest. The dataset provides 5-minute traffic-flow measurements from 8,600 loop-detector sensors across California’s highway system, collected from 2017–2021. Our experiments focus on the year 2020.
We discard sensors with more than 50% missing values. For the remaining sensors, missing entries are linearly interpolated between observed measurements, with leading or trailing gaps filled using forward/backward propagation. Sensors are randomly partitioned into training, validation, and test splits using a 6:1:3 ratio.
Each sensor’s time series is segmented into non-overlapping 14-day windows (4,032 time steps), yielding 26 windows per sensor. Within each window, timestamps are reset to start at 0 and increment in 5-minute steps. We then downsample each window by a factor of 6, producing a 30-minute resolution. Time indices are rescaled to days by dividing by $(60\times 24)$. Traffic-flow values are normalized to the $[0,1]$ range using min–max statistics computed from the training set.
The resulting dataset contains:
@@ B.3.2 Training and Data Description
Models are trained for 100 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. Each epoch consists of approximately 4,160 iterations, each processing a batch of 32 tasks. For every batch, the numbers of context and query points are drawn independently as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$ and $|\mathcal{I}_{q}|\sim\mathcal{U}[5,25)$; the sampled values are shared across all tasks within the batch. A task is formed by independently sampling $|\mathcal{I}_{c}|+|\mathcal{I}_{q}|$ time steps from a window without replacement. Training tasks are sampled on-the-fly from training windows. For validation and testing, we pre-generate fixed sets of tasks:
@@ B.4.1 Model Architectures
As with the other models, all MLP hidden sizes are increased from $256$ to $512$. We adopt the Efficient Query TNP design described in Section B.1.1, which employs a two-branch structure in which context embeddings are processed twice and query embeddings once. In contrast to previous sections, we allocate separate parameters for the two branches: at each layer, the multi-head self-attention over contexts and the multi-head cross-attention between queries and contexts no longer share parameters. For each multi-head attention module, the feedforward width is increased from $256$ to $512$, and all input/output embedding dimensions are set to $512$. All other configurations follow Section B.1.1.
@@
Because the data lie on a regular grid, we use the on-the-grid implementation of ConvCNP [gordon2019convolutional], eliminating the discretization step and improving computational efficiency. Let $\mathrm{I}$ denote an incomplete image, with unobserved pixels filled with dummy values, and let $\mathrm{M}_{c}$ be the binary mask indicating observed (context) pixels. For multi-channel images, the mask is broadcast along the channel dimension.
As in Section B.1.1, the output of the convolutional deep set module has two components: (i) a density channel capturing the spatial distribution of context pixels, and (ii) a kernel-smoothed representation. The kernel is implemented via a 2D convolutional layer (kernel size 11, $d_{{\mathbb{Y}}}$ input channels, 128 output channels, no bias) with a nonnegativity constraint enforced by taking absolute values of the learned weights during the forward pass [gordon2019convolutional].
@@
The SConvCNP mirrors the on-grid ConvCNP architecture described above, with the distinction that the ResNet backbone is replaced by a U-shaped FNO. To construct positional information, we uniformly discretize the interval [1,1][-1,1] along each spatial axis into grids whose sizes match the corresponding image dimensions. Their Cartesian product yields the 2D positional encoding, which is concatenated with the output of the convolutional deep set. This concatenated tensor is then passed pointwise through an MLP with one hidden layer of width 128 and GELU activations.
@@
B.4.2 Data Description
For our image-completion experiments, we use the Describable Textures Dataset (DTD; cimpoi2014describing). We adopt the standard DTD split, which contains 1,880 images each for training, validation, and testing. All images are RGB, with heights ranging from 231 to 778 pixels and widths from 271 to 900 pixels. During training, validation, and testing, we sample from each image a random $192\times 192$ crop, which we then resize to $64\times 64$. Pixel coordinates of the resulting $64\times 64$ grid along each axis are linearly mapped to [1,1][-1,1], and pixel intensities are normalized to $[0,1]$ independently across channels.
@@ B.5 Ablation Studies
We performed a series of ablation studies to quantify the contribution of three key design choices in SConvCNP: (i) the number of retained Fourier modes, (ii) the discretization resolution of the functional embedding, and (iii) the inclusion of positional encodings. Experiments were conducted on two 1D function families—samples from a Gaussian process with a Matérn–5/2 kernel and a sawtooth wave. To isolate the effect of each component, we used a slightly reduced-capacity SConvCNP model (3.1M parameters, 32 Fourier modes). Training followed the same protocol as in Section B.1, and evaluation was performed on 8,096 test tasks. A summary of our findings is provided below.
@@
Increasing the number of Fourier modes substantially improves performance on the sawtooth signal but yields negligible gains for Matérn–5/2 functions (Table 10). This difference is spectral: the sawtooth’s Fourier amplitudes decay only as $1/\xi$, leaving substantial energy at high frequencies and requiring many modes for accurate representation. In contrast, Matérn–5/2 samples exhibit spectra decaying as $1/\xi^{6}$, with extremely weak high-frequency content; thus, a small number of low-frequency modes already captures nearly all signal energy.
Table 10: Predictive log-likelihood of SConvCNP for different numbers of Fourier modes $m$ (4 seeds).
@@
Finer discretization improves performance in the sawtooth case but has minimal effect for Matérn–5/2 samples (Table 11). The sawtooth’s slowly decaying spectrum implies a high Nyquist frequency, so denser sampling is required to resolve its high-frequency components. In contrast, the Matérn–5/2 spectrum concentrates almost all energy at low frequencies; coarser discretizations already capture the relevant structure.
Table 11: Predictive log-likelihood of SConvCNP for different discretization resolutions (number of points per unit; 4 seeds).
@@
Adding positional encodings yields consistent improvements, particularly for the sawtooth wave (Table 12).
Table 12: Predictive log-likelihood of SConvCNP with and without positional encoding (4 seeds).

### v5

@@ B.3.1 Model Architectures
The model architectures follow those described in Section B.1.1. For the discretization of the functional embedding in both the ConvCNP and the SConvCNP, we use an expanded margin of 0.5 and a resolution of 48 grid points per unit length.
B.3.2 Data and Experimental Setup
We use the California traffic-flow dataset from LargeST [liu2023largest], a large-scale benchmark for traffic forecasting, available at https://www.kaggle.com/datasets/liuxu77/largest. The dataset contains traffic-flow measurements from 8,600 loop-detector sensors deployed across California’s highway network, collected at 5-minute intervals between 2017 and 2021. In our experiments, we restrict attention to data from the year 2020.
As a preprocessing step, we discard sensors with more than 50% missing values. For the remaining sensors, missing entries are filled via linear interpolation between observed measurements, with leading and trailing gaps completed using forward and backward propagation, respectively. Sensors are then randomly partitioned into training, validation, and test sets using a 6:1:3 split.
Each sensor’s time series is segmented into non-overlapping, continuous 14-day windows (4,032 time steps), yielding 26 windows per sensor. Within each window, timestamps are reset to start at 0 and increase in 5-minute increments. We subsequently downsample each window by a factor of 6, resulting in a temporal resolution of 30 minutes. Time indices are rescaled to units of days by dividing by $(60\times 24)$. Finally, traffic-flow values are normalized to the $[0,1]$ range using min–max statistics computed from the training set.
The processed dataset contains:
@@ B.3.2 Data and Experimental Setup
Models are trained for 100 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. Each epoch consists of approximately 4,160 iterations, each processing a batch of 32 tasks. For every batch, the numbers of context and query points are drawn independently as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$ and $|\mathcal{I}_{q}|\sim\mathcal{U}[5,25)$; the sampled values are shared across all tasks within the batch. A task is formed by independently and uniformly sampling $|\mathcal{I}_{c}|+|\mathcal{I}_{q}|$ time steps from a window without replacement. Training tasks are sampled on-the-fly from training windows. For validation and testing, we pre-generate fixed sets of tasks:
@@ B.4.1 Model Architectures
As with the other models, all MLP hidden sizes are increased from $256$ to $512$. We adopt the Efficient Query TNP design described in Section B.1.1, which employs a two-branch structure in which context embeddings are processed twice and query embeddings once. In contrast to previous sections, we allocate separate parameters for each branch: at each layer, the multi-head self-attention over contexts and the multi-head cross-attention between queries and contexts no longer share parameters. For each multi-head attention module, the feedforward width is increased from $256$ to $512$, and all input/output embedding dimensions are set to $512$. All other configurations follow Section B.1.1.
@@
Because the data lie on a regular grid, we use the on-the-grid implementation of the ConvCNP [gordon2019convolutional], eliminating the discretization step and improving computational efficiency. Let $\mathrm{I}$ denote an incomplete image, with unobserved pixels filled with dummy values, and let $\mathrm{M}_{c}$ be the binary mask indicating observed (context) pixels. For multi-channel images, the mask is broadcast along the channel dimension.
As in Section B.1.1, the output of the convolutional deep set module has two components: (i) a density channel capturing the spatial distribution of context pixels, and (ii) a kernel-smoothed representation. The kernel is implemented via a 2D convolutional layer (kernel size 11, $d_{y}$ input channels, 128 output channels, no bias) with a nonnegativity constraint enforced by taking absolute values of the learned weights during the forward pass [gordon2019convolutional].
@@
The SConvCNP mirrors the on-the-grid ConvCNP architecture described above, with the distinction that the ResNet backbone is replaced by a U-shaped FNO. To construct positional information, we uniformly discretize the interval [1,1][-1,1] along each spatial axis into grids whose sizes match the corresponding image dimensions. Their Cartesian product yields the 2D positional encoding, which is concatenated with the output of the convolutional deep set. This concatenated tensor is then passed pointwise through an MLP with one hidden layer of width 128 and GELU activations.
@@
B.4.2 Data and Experimental Setup
For our image-completion experiments, we use the Describable Textures Dataset (DTD; cimpoi2014describing). We adopt the standard DTD split, which contains 1,880 images each for training, validation, and testing. All images are RGB, with heights ranging from 231 to 778 pixels and widths from 271 to 900 pixels. During training, validation, and testing, we sample from each image a random $192\times 192$ crop, which we then downsample to $64\times 64$. Pixel coordinates of the resulting $64\times 64$ grid along each axis are linearly mapped to [1,1][-1,1], and pixel intensities are normalized to $[0,1]$ independently across channels.
@@ B.5 Ablation Studies
We conduct a series of ablation studies to assess the contribution of three core design choices in the SConvCNP:
  1. 1.
    the number of retained Fourier modes,
  2. 2.
    the discretization resolution of the functional embedding, and
  3. 3.
    the use of positional encodings.
Experiments are conducted on two families of one-dimensional functions: samples drawn from a GP with a Matérn–$5/2$ kernel, and a deterministic sawtooth waveform. Training follows the protocol described in Section B.1, with the only difference being the use of a smaller SConvCNP model (3.1M parameters, 32 Fourier modes). Results are reported over 8,096 test tasks, and the main findings are summarized below.
@@
Table 10: Predictive log-likelihood of SConvCNP for different numbers of Fourier modes $m$.
@@
Increasing the number of retained Fourier modes yields substantial improvements for the sawtooth signal, while having a negligible effect on Matérn–$5/2$ functions (Table 10). This behavior aligns with the spectral properties of the underlying functions. The sawtooth wave exhibits Fourier amplitudes that decay as $1/\xi$, leaving significant energy at high frequencies and necessitating a large number of modes for accurate reconstruction. In contrast, Matérn–$5/2$ samples have power spectra that decay as $1/\xi^{6}$, resulting in extremely weak high-frequency content; consequently, a small set of low-frequency modes suffices to capture nearly all signal energy.
@@
Table 11: Predictive log-likelihood of SConvCNP for different discretization resolutions (number of points per unit).
@@
A finer discretization improves performance for the sawtooth signal but has little impact on Matérn–$5/2$ samples (Table 11). The slowly decaying spectrum of the sawtooth implies a high effective Nyquist frequency, requiring dense sampling to resolve sharp transitions. By contrast, the spectral mass of Matérn–$5/2$ functions is concentrated at lower frequencies, so coarser discretizations can adequately capture the relevant structure.
@@
Table 12: Predictive log-likelihood of SConvCNP with and without positional encoding.
@@
Incorporating positional encodings consistently improves predictive performance, with particularly pronounced gains on the sawtooth tasks (Table 12). This suggests that explicit location information helps disambiguate high-frequency and non-smooth patterns that are not fully captured by fully translation equivariant features alone in the SConvCNP.